55 drops of oil each of volume 0.036cm cubic were released from a burette which was originally reading 10cm cubic .find the final reading of the burette
1
Expert's answer
2020-08-06T16:46:10-0400

The total volume of oil released from a burette will be:


"V = number\space of \space drops \times volume \space of \space drop = 55\times 0.036 cm^3 = 1.98cm^3"

Thus, the new reading will be:


"10cm^3 - V = 10cm^3 - 1.98cm^3 =8.02cm^3"

Answer. 8.02 cm cubic.